OpenStudy (anonymous):

Let f (x) = 3x – 6 and g(x) = x –2. Find f/g and its domain.

OpenStudy (anonymous):

I think f(x)=3x-6 g(x)=x-2 f/g =(3x-6)/(x-2) so the domain would be x≠2 Also you might get more (or better answers) if you put this in geometry next time :)